OpenStudy (anonymous):

sqrt 50 over 3 sqrt 15

OpenStudy (ghazi):

\[\frac{ \sqrt 50 }{ 3\sqrt {15} }=\]

OpenStudy (zzr0ck3r):

sqrt(50) = sqrt(25*2) = sqrt(25)*sqrt(2) = 5sqrt(2) hope this helps

OpenStudy (anonymous):

yeah ghazi

OpenStudy (ghazi):

\[\frac{ 5 \sqrt 2 }{ 3 \sqrt 15 }*\frac{ \sqrt 15 }{ \sqrt15 }= \frac{ 5 \sqrt{30} }{ 45 }\]

OpenStudy (zzr0ck3r):

touche

OpenStudy (anonymous):

i think the answer is 25 sqrt 30 over 45

OpenStudy (anonymous):

ghazi that is not a choice

OpenStudy (ghazi):

well solution is in front of you...

OpenStudy (anonymous):